OpenStudy (anonymous):

Which of the following is a solution to √x-4+9=6 this would be -7 right thanks btw.

OpenStudy (mertsj):

\[\sqrt{x-4}+9=6\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yes sir.

OpenStudy (mertsj):

Did you subtract 9 from both sides?

OpenStudy (mertsj):

\[\sqrt{x-4}=-3\]

OpenStudy (anonymous):

then +4

OpenStudy (anonymous):

it would just be 1

OpenStudy (anonymous):

x=1

OpenStudy (mertsj):

\[\sqrt{x+4}=-3\]

OpenStudy (anonymous):

oh -7.

OpenStudy (mertsj):

Are you aware that this symbol: \[\sqrt{}\]

OpenStudy (anonymous):

what does that mean then?

OpenStudy (mertsj):

means the POSITIVE square root?

OpenStudy (anonymous):

it is 1 then.

OpenStudy (mertsj):

So you have a statement that says that the POSITIVE square root is equal to NEGATIVE 3.

OpenStudy (mertsj):

Can a positive number be equal to a negative number?

OpenStudy (anonymous):

no.

OpenStudy (mertsj):

Then what would your conclusion be?

OpenStudy (anonymous):

its -3

OpenStudy (mertsj):

Look at me: Can a positive number equal a negative number?

OpenStudy (anonymous):

no

OpenStudy (mertsj):

Supposing I say to you: Find a positive number x such that x = -9. What would you say.

OpenStudy (anonymous):

9.

OpenStudy (mertsj):

Ok. Let me tell you what you should say. You should say, "There is no such number because a positive number cannot equal a negative number."

OpenStudy (anonymous):

THE ANSWER IS 1.

OpenStudy (mertsj):

So let me get this straight:

OpenStudy (anonymous):

I dont have all day i wanna go to bed hurry up. Whats the answer.

OpenStudy (mertsj):

\[\sqrt{1-4}+9=6\]

OpenStudy (anonymous):

THE ANSWER IS 13.

OpenStudy (mertsj):

The answer is no solution. And stop with the foul language.

OpenStudy (anonymous):

oh really because there is no option that says no solution.

OpenStudy (mertsj):

What is so hard to understand that there is no number that makes the sentence true.

OpenStudy (mertsj):

Then you haven't posted the problem correctly.

OpenStudy (anonymous):

x = -3 x = -7 x = 1 x = 13

OpenStudy (anonymous):

go to the top read what I posted there is not a 1 for x.

OpenStudy (anonymous):

its